OpenStudy (anonymous):

A liquid is filled up to a level of (h) in a water tank of area (A). This tank is taken up with an acceleration of (a). Find the net force acting on the bottom of the tank. ( density of the liquid=d, acceleration due to gravity=g ) please help.

OpenStudy (anonymous):

please i need help :(

OpenStudy (anonymous):

mass of liquid = density of liquid * volume of liquid = d * hA =dhA Now, Net force = mass of liquid * acceleration = dhA *a =dhAa

OpenStudy (anonymous):

got it?

OpenStudy (agent0smith):

@sauravshakya i don't think that's completely right. The net force is still given by the weight, mg, and that force is increasing because the weight is increasing as water is pumped in (i think?) "This tank is taken up with an acceleration of (a)" I'm not sure if this means the tank is being lifted up with an acceleration a, or that the tank is being filled at a rate a... i think it means it's being lifted up with an acceleration a, in which case your formula just needs the addition of g: Net force = mass of liquid * ( gravity + acceleration) = dhA *(g + a) =dhA(g+a)

OpenStudy (anonymous):

I guess the tank is lifted up with an acceleration 'a' So, NET FORCE =mass * acceleration

OpenStudy (shubhamsrg):

i'd agree with @agent0smith , apart from the pseudo force ma, mg is always acting on it, so net accn will be g+a .

OpenStudy (anonymous):

In this case find the NET FORCE